OpenStudy (anonymous):

When you solve an equation, how do you know when to add and when to subtract? You may include an example if you need to.

OpenStudy (anonymous):

what type of equation?

OpenStudy (anonymous):

what math are you in?

OpenStudy (anonymous):

Solving Equations Using Addition. As in example: x minus forty-five equals twelve

OpenStudy (anonymous):

so... x-54=12 ; you will add 54 to the other side because you wanna get rid of it, but when you do that you have o make sure you add it on the other side too so it will look like: x-54=12 +54 +54 now there is only an x on one side and on the other side its now 66. (its 66 because you added 54)

OpenStudy (anonymous):

so it looks like this x=66, and that is your awnser

OpenStudy (anonymous):

Yes, I get that. Yet I don't need to solve anything, I just need to explain 'when I know to subtract' (:

OpenStudy (anonymous):

now for your question, see how I added the 54? I added because the 54 is a negative in the equation.

OpenStudy (anonymous):

you subtract when there is a positive and you add when there is a negative. if it was x+54, then you would subtract. but seince its -54 you add

OpenStudy (anonymous):

the point of subtracting is to get rid of the number on that side so the x can be all alone